Derf Corporation uses a standard cost system in which it applies manufacturing overhead on the basis of standard direct labor-hours. Two direct labor-hours are required for each unit produced. The denominator activity was set at 6,400 units. Manufacturing overhead was budgeted at $102,400 for the period; 20 percent of this cost was fixed. The 11,670 hours worked during the period resulted in production of 5,710 units. Variable manufacturing overhead cost incurred was $82,420 and fixed manufacturing overhead cost was $22,230.

The variable overhead rate variance for the period was:
Answer #1

Standard rate = 102400/12800 = 8*80% = 6.4

Variable overhead rate variance = (Standard rate-actual rate)actual hour

= (6.4*11670-82420)

Variable overhead rate variance = 7732 U